# A multiverse model in dS wedge holography

### v2

@@ Abstract
We construct a multiverse model based on a recent proposal for dS wedge holography, where empty AdS3 space is cut off by a pair of accelerated dS2 space universes, one near the AdS boundary, denoted the UV brane, and one in the AdS interior, the IR brane. We glue together several copies of this configuration along the UV and IR branes in a periodic matter. To provide the model with dynamics similar to those of near Nariai black holes used in other multiverse toy models, we add dS JT gravity as an intrinsic gravity theory on the IR branes. We then study the entanglement entropy with respect to a UV brane observer, who finds a Page curve transition due to an entanglement island connecting the UV and IR branes. This process involves the coarse-graining of information outside the causally accessible region to the observer. Our model provides an explicit realization of entanglement between IR and UV degrees of freedom encoded in the multiverse.
@@ 1 Introduction
Since the ETW branes in wedge holography can have arbitrary cosmological constants, this has led to different realizations of wedge holography, and importantly for us, to the development of de Sitter (dS) wedge holography Aguilar-Gutierrez:2023tic . Here, one employs a dSd space ETW branes near the asymptotic boundary of AdSd+1 space and in the interior bulk geometry, denoted as the ultraviolet (UV) and infrared (IR) branes respectively. [^1] There are also notions of flat/dS space wedge holography; see Ogawa:2022fhy ; Bhattacharjee:2022pcb for recent developments. Using this setting, one can study different notions of quantum information with respect to an observer living in the UV brane universe, since gravity is not dynamical in this region. [^2] Other approaches to quantum information on dS braneworlds can be found in Geng:2021wcq ; Yadav:2023qfg ; Cotler:2022weg . In particular, for AdS3 ambient space, one can analytically reproduce a Page curve with respect to a UV brane observer. Given that pure Einstein gravity is topological in 3-dimensions, one normally either introduces fluctuations in the braneworld location Geng:2022slq ; Geng:2022tfc , or intrinsic gravity theories on the brane. The latter approach has been developed in different areas, such as for providing new hints in the context of the information paradox Chen:2020uac ; Chen:2020hmv ; Miao:2023unv ; Li:2023fly , and developing bounds the intrinsic gravity couplings on the ETW branes based on consistency with entanglement velocity Lee:2022efh . The higher curvature corrections to the intrinsic gravity on the brane render the gravity dynamical, as opposed to purely topological. This allows for a more consistent holographic treatment of branes with intrinsic gravity and sharp formulations of quantum information observables Lee:2022efh . It has been found that introducing the intrinsic gravitational theory reproduces a Page curve with massless gravitons in AdS wedge holography Miao:2023unv , and cone holography Li:2023fly .
@@
Our work aims at exploring the coarse-graining of information encountered in semi-classical quantum cosmology Hartle:2016tpo ; Aguilar-Gutierrez:2021bns within dS wedge holography. In the context of eternal inflation in quantum cosmology, such as false vacuum eternal inflation [^3] False vacuum decay refers to a tunneling event from a local minimum in an inflaton potential to global minimum. The reader is refereed to Coleman:1980aw for a pionnering publication in vacuum decay in semiclassical gravity, and Garriga:2006hw for original work on false vacuum eternal inflation., one expects that meta-observers in a given universe have access to the information available in other universes given a large amount of redundancy in the theory Hartle:2016tpo ; Aguilar-Gutierrez:2021bns , and yet they cannot interact with those (spacelike separated) universes. This coarse-grained information shares similarities with the fine-grained von Neumann entropy computed with entanglement islands Penington:2019npb ; Almheiri:2019psf , as both are determined from the saddle points from the semi-classical gravitational path integral i.e. a coarse-graining over the allowed geometrical configurations of the theory.
@@
In this work, we study this situation within a multiverse dS-braneworld toy model, to gain access to holographic tools that allow us to evaluate the holographic entanglement entropy of a subregion in a multiverse through the Hubeny-Rangamani-Takayanagi (HRT) formula Ryu:2006bv ; Ryu:2006ef ; Hubeny:2007xt . This will help us interpret the information a UV brane observer has access to in terms of CFT degrees of freedom. To carry out this task, we introduce the braneworld model illustrated in Fig 1.444Throughout the letter, we employ latin indices for braneworld coordinates.
@@
Figure 1: “dS wedge holographic” multiverse model formed with a pair of dSd branes, labeled by Qb (in purple) and Qc (gray) embedded in AdSd+1 space (orange), and ending on a pair of codimension-two Euclidean defects at global AdS time $\tau=\pm\frac{\pi}{2}$. The region between the branes (blue) is denoted as the wedge Wd+1. We glue together different copies of the same system along the UV and IR branes (represented by the red arrows) to form a double-sided brane wedge configuration for both the UV and IR branes. From the brane perspective, there is a junction between dSd universes on opposite sides. The respective Penrose diagram is shown in Fig. 2.
@@
This model can be viewed as an extension of dS wedge holography Aguilar-Gutierrez:2023tic . A technical difference with Aguilar-Gutierrez:2023tic is that besides connecting different bulk spacetimes through the dS branes, we also add an intrinsic gravity theory in the IR ETW branes. This allows for the effective braneworld description to contain dynamics even for ($2+1$)-dimensional bulk space with dS2 space branes. We adopt dS Jackiw-Teitelboim (JT) gravity JACKIW1985343 ; TEITELBOIM198341  [^5] This theory can describe the perturbations around the near horizon region of arbitrary dimensional near extremal Schwarzschild-de Sitter (SdS) black holes, known as the near Nariai limit, or from the dimensional reduction of pure dS3 space. We will consider only the near Nariai perspective. See, for instance, Svesko:2022txo and references therein for general aspects of these theories. as the intrinsic theory on the IR branes, while keeping the UV branes to be pure tensional. This is achieved by imposing Dirichlet (Neumann) boundary conditions on the UV (IR) brane, and it allows us to fix the cosmological constant in both the UV and IR branes to be dS space. Having intrinsic gravity on the brane modifies the gravitational observables anchored to the UV boundary, such that we can study the consequences for the properties in the multiverse model.
@@
Our main point of interest is the evolution of holographic entanglement entropy with respect to the CFT degrees of freedom in the UV branes within this model. [^6] See Levine:2022wos ; Yadav:2023qfg ; Aguilar-Gutierrez:2021bns ; Pasquarella:2022ibb ; Jiang:2024xnd ; Yadav:2024ray for closely related previous studies. This construction has an apparent black hole-like information loss paradox. Based on Aguilar-Gutierrez:2023tic , we consider a UV brane observer collecting Hawking radiation. [^7] Having the brane very close to the asymptotic boundary allows for suppressing higher curvature corrections on the brane. It was shown in Iwashita:2006zj that the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef is generically not applicable in a dS braneworld embedded in higher dimensional AdS space due to the presence of higher derivative corrections. If instead, we place the brane at the conformal boundary where these corrections are completely suppressed, the brane theory decouples from the bulk gravity, also preventing any non-trivial interaction between the two. The von Neumann entropy with respect to the UV observer describes two phases; at early times, it is a monotonically increasing function of time during the so-called Hawking phase. The UV observer could access more Hawking modes than the total number of degrees of freedom on the IR brane union with the entangling region in question. However, later there is a phase where it decreases and then saturates, which is an example of the entanglement islands Penington:2019npb ; Almheiri:2019psf .
@@
Our results suggest that the information available to each observer living in a given UV brane completely encodes that of the other universes, as a consequence of the Israel junction conditions Israel:1966rt in the configuration (see Sec. 3). This observation confirms that the braneworld multiverse model explicitly realizes the coarse-graining of information for observables in quantum cosmology Hartle:2016tpo ; Aguilar-Gutierrez:2021bns . Moreover, the island configurations we encounter show that the different observers cannot transmit nor decode messages between universes. This result is consistent with the central dogma Shaghoulian:2021cef and the no-cloning theorem for cosmological horizons discussed in Levine:2022wos . The central dogma in this context refers to encoding information of the spacetime beyond the cosmological horizon, with respect to a given observer, from the interior region. In Levine:2022wos , it was noticed in a multiverse toy model that observers in spacelike separated regions could in principle encode regions of spacetime with some overlap, such that they could reconstruct information from those regions without affecting the ability of the other observer to do the same. This would then enter into tension with the no-cloning theorem. However, the authors also find additional non-overlapping island saddles with respect to meta-observers on a non-dynamical gravity region. The dominance of these saddles implies that the apparent paradox simply does not arise. The model in Levine:2022wos has many similarities to ours. The non-dynamical region in our setting refers instead to the UV brane where one collects Hawking radiation. The island transitions show that the RT surfaces will be confined to a single dS wedge universe. The main new observation, explicit in our model, is the entanglement between the UV and IR degrees of freedom present in the multiverse model, and in the coarse-graining of information with respect to different UV brane observers, who take the role of the meta-observers in the false vacuum eternal inflation models.
@@
Outline: The rest of this paper is organized as follows. We start by reviewing dS wedge holography and we introduce our multiverse braneworld model in section 2. In section 3, we present an information recovery protocol and compute the holographic entanglement entropy of spacetime subregions using the HRT formula, which results in a page curve transition with respect to a UV braneworld observer. In section 3.5 we comment on the connection between our model with the central dogma and non-cloning theorem in the context of quantum cosmology. We conclude in section 4 with a discussion of our findings and important questions to be addressed in the future.
@@ 2 Braneworld (multiverse) model
To formulate the multiverse model, we start constructing its building block as a single AdSd+1 bulk geometry with two ETW branes, one inside the bulk, denoted the IR brane, and the other near the asymptotic boundary, the UV brane. This simple model first appeared in Aguilar-Gutierrez:2023tic as a new proposal for dS wedge holography. In this section, we first review the basic construction involving the single UV and IR brane configuration without intrinsic gravity, and later, we introduce multiple ETW branes with an intrinsic dS JT gravity theory on them.
2.1 Review of dS wedge holography
The original proposal in dS wedge holography Aguilar-Gutierrez:2023tic considered a double holographic AdSd+1 space bounded by a single pair of dSd space near the asymptotic boundary (denoted as the UV region, where gravity decouples) and an arbitrary location in the bulk interior (the IR region). This brane configuration in the Hartle-Hawking preparation of state corresponds to a tunneling instanton describing membrane creation Aguilar-Gutierrez:2023tic .
@@ 2.1 Review of dS wedge holography
    AdSd+1 bulk space with a pair of dSd Randall-Sundrum branes Randall:1999ee  [^8] As noticed in an alternative Randall-Sundrum multiverse model Yadav:2023qfg , the dS branes would have a finite lifetime. In our configuration the accelerating branes produce a singularity at the location where the branes nucleate Garriga:1993fh ; Arcos:2022icf (a big bang), as well as where they decay (a big crunch) Emparan:2022ijy , where the lack of unitary on the codimension-two dS conformal defects is manifest. that overlap at global AdS time $\tau=\pm\frac{\pi}{2}$.
@@
In the above $h_{\mu\nu}$ is the induced metric on $\partial\mathcal{M}$; ^$\hat{\mathcal{L}}_{\rm bulk}$ is the bulk field Lagrangian density; $Q_{c}$ and $Q_{b}$ denote (the worldvolume) of the UV and IR ETW branes (which we consider to be double sided) respective; $\mathcal{L}^{(i)}_{\rm intrinsic}$ is the intrinsic gravity theory on the $Q_{i}$ brane; $h^{(i)}_{ij}$ the respective induced metric (where $i,\,j$ are $2$-dimensional indices); $I^{(i)}_{\rm matter}$ is the matter field theory on the branes; and $G_{d+1}$ is Newton’s gravitational constant.
@@
In this foliation, dS branes of arbitrary tension appear can only be found in the range $-\frac{\pi}{2}\leq\tau\leq\frac{\pi}{2}$ Parikh:2012kg . [^9] The UV and IR ETW branes overlap at global AdS time $\tau=\pm\pi/2$ as seen from (6) for $t\rightarrow\pm\infty$ and $\rho\rightarrow\infty$ in both branes.
We can also employ a change of coordinates from global AdSd+1 space to AdSd+1 with dSd space foliations to place the UV ETW branes, which will have an effective positive cosmological. In general, the metric has the form Arcos:2022icf ,
@@
where $\mathrm{d}s_{\rm dS}$ is a $d$-dimensional line element for dS space in any coordinate system, and $H$ is the Hubble rate. In these coordinates, we locate the IR and UV ETW branes at $\sigma=\sigma_{b}$ and $\sigma=\sigma_{c}$ respectively, such that
@@
To describe the evolution with respect to the global time of an observer living in the dS ETW brane, it is most convenient to use a Rindler-AdSd+1 background with global coordinate dSd foliation with the explicit mapping

### v3

# A multiverse model in $T^{2}$ dS wedge holography

@@ Abstract
We construct a multiverse model where empty AdSd+1 space is cut off by a pair of accelerated dSd space universes, at a finite AdS boundary cutoff which we treat as a $T^{2}$ deformation in the holographic dual, and one in the AdS interior, the IR brane; and denote the construction as $T^{2}$ dS wedge holography. We glue together several copies of this configuration along the UV cutoff and the IR branes in a periodic matter. To provide the model with dynamics similar to those of near Nariai black holes used in other multiverse toy models, we specialize to $d=2$ and add dS JT gravity as an intrinsic gravity theory on the IR branes. We then study the entanglement entropy with respect to a finite cutoff observer, who finds a Page curve transition due to an entanglement island connecting the UV cutoff and IR brane. This process involves the coarse-graining of information outside the causally accessible region to the observer. Our model provides an explicit realization of entanglement between IR and UV degrees of freedom encoded in the multiverse.
@@ 1 Introduction
Since the ETW branes can have arbitrary cosmological constants, this has led to different realizations of wedge holography, and importantly for us, to the development of de Sitter (dS) wedge holography Aguilar-Gutierrez:2023tic . Here, one employs a dSd space ETW branes near the asymptotic boundary of AdSd+1 space and in the interior bulk geometry, denoted as the ultraviolet (UV) and infrared (IR) branes respectively. [^1] There are also notions of flat/dS space wedge holography; see Ogawa:2022fhy ; Bhattacharjee:2022pcb for recent developments. By placing the UV brane sufficiently close to the boundary to effectively render the gravity non-dynamical, Aguilar-Gutierrez:2023tic studied the holographic entanglement entropy Ryu:2006bv ; Ryu:2006ef ; Hubeny:2007xt ; Faulkner:2013ana ; Engelhardt:2014gca and holographic complexity Susskind:2014rva ; Stanford:2014jda ; Belin:2021bga ; Belin:2022xmt with respect to an observer living in the UV cutoff universe to be studied [^2] Other approaches to quantum information on dS braneworlds can be found in Geng:2021wcq ; Yadav:2023qfg ; Cotler:2022weg .. In particular, for AdS3 ambient space, one can analytically reproduce a Page curve with respect to a UV observer. Given that pure Einstein gravity is topological in 3-dimensions, one normally either introduces fluctuations in the braneworld location Geng:2022slq ; Geng:2022tfc , or intrinsic gravity theories on the brane. The latter approach has allowed progress in different areas, such as for providing new hints in the context of the information paradox Chen:2020uac ; Chen:2020hmv ; Miao:2023unv ; Li:2023fly , and developing bounds the intrinsic gravity couplings on the ETW branes based on consistency with entanglement velocity Lee:2022efh . The higher curvature corrections to the intrinsic gravity on the brane render the gravity dynamical, as opposed to purely topological. This allows for a more consistent holographic treatment of branes with intrinsic gravity and sharp formulations of quantum information observables Lee:2022efh . It has been found that introducing the intrinsic gravitational theory reproduces a Page curve with massless gravitons in AdS wedge holography Miao:2023unv , and cone holography Li:2023fly .
@@
Our work aims at exploring the coarse-graining of information encountered in semi-classical quantum cosmology Hartle:2016tpo ; Aguilar-Gutierrez:2021bns by proposing an extension to dS wedge holography, where instead of a second ETW brane near the asymptotic AdS boundary, we locate a finite boundary cutoff generated by a $T^{2}$ deformation Hartman:2018tkw in the dual CFT, which is kept very close to the asymptotic AdS boundary - we refer to this as $T^{2}$ dS wedge holography. We perform this step in order to impose Dirichlet boundary conditions at a finite radial location333 Different works of wedge holography with non-trivial observables consider that at least one of the ETW branes is connected with a non-gravitating bath, such as a boundary CFT Caceres:2021fuw ; Geng:2021mic ; Ghosh:2021axl ; Caceres:2020jcn ; Anderson:2020vwi ; Bhattacharya:2021jrn ; Geng:2021iyq ; Qi:2021sxb . Here, we will instead treat the regulated UV brane as a $T^{2}$ deformation, with a well-defined Dirichlet boundary at a finite fixed radial distance Zamolodchikov:2004ce ; Taylor:2018xcy ; Hartman:2018tkw , and evaluate quasi-local diffeomorphism invariant observables; where the closeness of the finite cutoff region to the asymptotic boundary allows us to neglect the effect of non-localities, at scales of the deformation parameter Cardy:2019qao , in the dual QFT Apolo:2023ckr ; Banerjee:2024wtl ; Demise:2021cfx ; Chen:2018eqk ; Donnelly:2018bef . In the context of eternal inflation in quantum cosmology, such as false vacuum eternal inflation [^4] False vacuum decay refers to a tunneling event from a local minimum in an inflaton potential to the global minimum. The reader is referred to Coleman:1980aw for a pioneering publication in vacuum decay in semiclassical gravity, and Garriga:2006hw for original work on false vacuum eternal inflation., one expects that meta-observers in a given universe have access to the information available in other universes given a large amount of redundancy in the theory Hartle:2016tpo ; Aguilar-Gutierrez:2021bns , and yet they cannot interact with those (spacelike separated) universes. This coarse-grained information shares similarities with the fine-grained von Neumann entropy computed with entanglement islands Penington:2019npb ; Almheiri:2019psf , as both are determined from the saddle points from the semi-classical gravitational path integral i.e. a coarse-graining over the allowed geometrical configurations of the theory.
@@
In this work, we study this situation within a multiverse dS-braneworld toy model, to gain access to holographic tools that allow us to evaluate the holographic entanglement entropy of a subregion in a multiverse through the Hubeny-Rangamani-Takayanagi (HRT) formula Ryu:2006bv ; Ryu:2006ef ; Hubeny:2007xt . This will help us interpret the information a UV observer has access to in terms of CFT degrees of freedom. To carry out this task, we introduce the braneworld model illustrated in Fig 1.555Throughout the letter, we employ latin indices for braneworld coordinates.
@@
Figure 1: “dS wedge holographic” multiverse model formed with a pair of dSd branes, labeled by Qb (in purple) and Qc (gray) embedded in AdSd+1 space (orange), and ending on a pair of codimension-two Euclidean defects at global AdS time $\tau=\pm\frac{\pi}{2}$. The region between the branes (blue) is denoted as the wedge Wd+1. We glue together different copies of the same system along the UV cutoff and IR brane (represented by the red arrows) to form a double-sided brane wedge configuration for both the UV cutoff and IR brane. From the brane perspective, there is a junction between dSd universes on opposite sides. The respective Penrose diagram is shown in Fig. 2.
@@
This model can be viewed as an extension of dS wedge holography Aguilar-Gutierrez:2023tic . A technical difference with Aguilar-Gutierrez:2023tic is that besides connecting different bulk spacetimes through the dS branes, we also add an intrinsic gravity theory in the IR ETW branes; besides generating the UV cutoff through a $T^{2}$ deformation in the dual CFT, as mentioned above. This allows for the effective braneworld description to contain dynamics even for ($2+1$)-dimensional bulk space with a dS2 space brane in the IR region of the bulk and a UV cutoff. We adopt dS Jackiw-Teitelboim (JT) gravity JACKIW1985343 ; TEITELBOIM198341  [^6] This theory can describe the perturbations around the near horizon region of arbitrary dimensional near extremal Schwarzschild-de Sitter (SdS) black holes, known as the near Nariai limit, or from the dimensional reduction of pure dS3 space. We will consider only the near Nariai perspective. See, for instance, Svesko:2022txo and references therein for general aspects of these theories. as the intrinsic theory on the IR branes, while keeping the UV cutoff. This is achieved by imposing Dirichlet (Neumann) boundary conditions on the UV (IR), and it allows us to fix the cosmological constant in both the finite cutoff region and IR brane to be dS space. Having intrinsic gravity on the brane modifies the gravitational observables anchored to the UV boundary, such that we can study the consequences for the properties in the multiverse model.
@@
Our main point of interest is the evolution of holographic entanglement entropy with respect to the $T^{2}$-deformed CFT degrees of freedom in the UV cutoff within this model. [^7] See Levine:2022wos ; Yadav:2023qfg ; Aguilar-Gutierrez:2021bns ; Pasquarella:2022ibb ; Jiang:2024xnd ; Yadav:2024ray for closely related previous studies. On the technical level, in order to have well-defined notion of holographic entanglement entropy, we need to restrict the analysis to a quasi-local description of the dual field theory (see e.g. Apolo:2023ckr ; Banerjee:2024wtl ), which is realized when the length of the entangling region in the HRT formula is much larger than the length scale set by the deformation parameter. Under this restriction for the entangling region, our construction has an apparent black hole-like information loss paradox. Based on Aguilar-Gutierrez:2023tic , we consider a UV observer collecting Hawking radiation. The von Neumann entropy with respect to the UV cutoff observer describes two phases; at early times, it is a monotonically increasing function of time during the so-called Hawking phase. The observer could access more Hawking modes than the total number of degrees of freedom on the IR brane union with the entangling region in question. However, later there is a phase where it decreases and then saturates, which is an example of the entanglement islands Penington:2019npb ; Almheiri:2019psf . In the connected phase, we find that the entanglement entropy is non-vanishing and depends on degrees of freedom in both the IR and the UV; the minimal surface after the Page transition is anchored to both of the branes.
@@
Our results suggest that the information available to each observer living in a given UV cutoff completely encodes that of the other universes, as a consequence of the Israel junction conditions Israel:1966rt in the configuration (see Sec. 3). This observation confirms that the braneworld multiverse model explicitly realizes the coarse-graining of information for observables in quantum cosmology Hartle:2016tpo ; Aguilar-Gutierrez:2021bns . Moreover, the island configurations we encounter show that the different observers cannot transmit nor decode messages between universes. This result is consistent with the central dogma Shaghoulian:2021cef and the no-cloning theorem for cosmological horizons discussed in Levine:2022wos . The central dogma in this context refers to encoding information of the spacetime beyond the cosmological horizon, with respect to a given observer, from the interior region. In Levine:2022wos , it was noticed in a multiverse toy model that observers in spacelike separated regions could in principle encode regions of spacetime with some overlap, such that they could reconstruct information from those regions without affecting the ability of the other observer to do the same. This would then enter into tension with the no-cloning theorem. However, the authors also find additional non-overlapping island saddles with respect to meta-observers on a non-dynamical gravity region. The dominance of these saddles implies that the apparent paradox simply does not arise. The model in Levine:2022wos has many similarities to ours. The non-dynamical region in our setting refers instead to the UV cutoff where one collects Hawking radiation. The island transitions show that the RT surfaces will be confined to a single dS wedge universe. The main new observation, explicit in our model, is the entanglement between the UV and IR degrees of freedom present in the multiverse model, and in the coarse-graining of information with respect to different UV observers, who take the role of the meta-observers in the false vacuum eternal inflation models.
Lastly, our proposal for a system with a UV and IR cutoff region also relates to recent works where double-trace $T\overline{T}$ deformations in a CFT dual to pure AdS3 space with an ETW brane allow for a new realization of the Page transition Deng:2023pjs . In contrast to this other approach, the dS cutoff regions do not overlap with each other, except at particular time slices (see Fig. 1) which modifies considerably the analysis, with consequences for the Page curve transition, and the quantum cosmology interpretation, which we persuade in Sec. 3.
@@
Outline: The rest of this paper is organized as follows. We start by reviewing the geometric construction and we introduce our multiverse braneworld model in Sec. 2. In Sec. 3, we present an information recovery protocol and compute the holographic entanglement entropy of spacetime subregions using the HRT formula, which results in a page curve transition with respect to a UV observer. In Sec. 3.5 we comment on the connection between our model with the central dogma and non-cloning theorem in the context of quantum cosmology. We conclude in Sec. 4 with a discussion of our findings and important questions to be addressed in the future.
@@ 2 Braneworld (multiverse) model
To formulate the multiverse model, we start constructing its building block as a single AdSd+1 bulk geometry with an ETW brane inside the bulk, denoted the IR brane, and the finite cutoff (produced by a $T^{2}$ deformation in the dual theory). This simple model is a modification of dS wedge holography that appeared in Aguilar-Gutierrez:2023tic . Importantly, instead of considering an ETW brane very close to the asymptotic boundary,we will consider a $T^{2}$ deformation of the original CFTd theory living on a dSd background dual to pure AdSd+1 with the appropriate foliation. This allows for consistency at the moment of imposing Dirichlet boundary conditions at a finite boundary location, which will allow us to find a Page curve transition in the multiverse toy model in Sec. 3.. In this section, we specify the geometric construction, including an IR brane without intrinsic gravity, and later, we introduce multiple ETW branes with an intrinsic dS JT gravity theory on them.
2.1 De Sitter wedge holography, and $T^{2}$ deformations
The original proposal in dS wedge holography Aguilar-Gutierrez:2023tic considered a double holographic AdSd+1 space bounded by a single pair of dSd space near the asymptotic boundary (denoted as the UV region, where gravity decouples) and an arbitrary location in the bulk interior (the IR region). The configuration was interpreted as the Hartle-Hawking preparation of state corresponds to a tunneling instanton describing membrane creation Aguilar-Gutierrez:2023tic in the IR. Our configuration, however, considers an important modification, where the finite cutoff UV region is instead due to a $T^{2}$ deformation in the dual CFT (see Sec. 2.2).
@@ 2.1 De Sitter wedge holography, and $T^{2}$ deformations
    AdSd+1 bulk space with a pair of dSd cutoff region, where the IR corresponds to a Randall-Sundrum brane Randall:1999ee  [^8] As noticed in an alternative Randall-Sundrum multiverse model Yadav:2023qfg , the dS branes would have a finite lifetime. In our configuration the accelerating branes produce a singularity at the location where the branes nucleate Garriga:1993fh ; Arcos:2022icf (a big bang), as well as where they decay (a big crunch) Emparan:2022ijy , where the lack of unitary on the codimension-two dS conformal defects is manifest. that overlap at global AdS time $\tau=\pm\frac{\pi}{2}$, while the UV region corresponds to a $T^{2}$ deformation Hartman:2018tkw in the dual CFTd.
@@
In the above $h_{\mu\nu}$ is the induced metric on $\partial\mathcal{M}$; ^$\hat{\mathcal{L}}_{\rm bulk}$ is the bulk field Lagrangian density; $Q_{c}$ is the UV cutoff region, and $Q_{b}$ denote (the worldvolume) of the IR ETW brane, where we consider both regions to be double sided; $\mathcal{L}^{(i)}_{\rm intrinsic}$ is the intrinsic gravity theory on the $Q_{i}$ regions; $h^{(i)}_{ij}$ the respective induced metric (where $i,\,j$ are $2$-dimensional indices); $I^{(b)}_{\rm matter}$ is the matter field theory on $Q_{i}$; and $G_{d+1}$ is Newton’s gravitational constant.
@@
In this foliation, dS branes of arbitrary tension can only be found in the range $-\frac{\pi}{2}\leq\tau\leq\frac{\pi}{2}$ Parikh:2012kg . [^9] The UV cutoff region and the IR ETW brane overlap at global AdS time $\tau=\pm\pi/2$ as seen from (6) for $t\rightarrow\pm\infty$ and $\rho\rightarrow\infty$ in both regions.
We can also employ a change of coordinates from global AdSd+1 space to AdSd+1 with dSd space foliations to place the ETW branes, which will have an effective positive cosmological. In general, the metric has the form Arcos:2022icf ,
@@
where $\mathrm{d}s_{\rm dS}$ is a $d$-dimensional line element for dS space in any coordinate system, and $H$ is the Hubble rate. In these coordinates, we locate the finite IR and UV cutoff regions at $\sigma=\sigma_{b}$ and $\sigma=\sigma_{c}$ respectively, such that
@@
To describe the evolution with respect to the global time of an observer living in the dS finite UV cutoff, it is most convenient to use a Rindler-AdSd+1 background with global coordinate dSd foliation with the explicit mapping
@@
2.2 Incorporating $T^{2}$ deformations
$T^{2}$ deformations were defined by Hartman:2018tkw  [^10] See also Taylor:2018xcy ; Morone:2024ffm . as a generalization of the finite cutoff interpretation of $\text{T}\overline{\text{T}}$ for CFTd in arbitrary dimensions. The definition is based on the $\text{AdS}_{d+1}/\text{CFT}_{d}$ dictionary relating the bulk gravity and CFT partition functions, which is taken to hold at finite bulk radial cutoff $r_{B}$
$Z_{\text{EFT}}[r_{B};\gamma_{ij},J]=Z_{\text{grav}}[h_{ij}^{B}=r_{B}^{2}\gamma_{ij},\psi_{B}=r_{B}^{\Delta-d}J]\;.$ (8)
The left-hand side is the generating function for the (assumed holographic) effective field theory, which need not be a CFT itself, $h_{ij}$ is the metric describing the field theory geometry, and $J$ is, for simplicity, taken to be a source for a scalar operator $\mathcal{O}$ of dimension $\Delta$. On the right-hand side is the (on-shell) gravitational partition function in an asymptotically AdS background with metric
$\mathrm{d}s^{2}=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}=\frac{\mathrm{d}r^{2}}{N(r)}+r^{2}\gamma_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}\leavevmode\nobreak\ ,$ (9)
where $N(r)\to r^{2}$ near the conformal boundary; the bulk metric and bulk scalar field $\psi$ are taken to obey Dirichlet boundary conditions, i.e. fixing the boundary-induced metric $h_{ij}(r_{B},x)\equiv h^{B}_{ij}(x)$; and bulk fields $\psi(r_{B},x)\equiv\psi_{B}(x)$. The standard dictionary is recovered in the limit $r_{B}\to\infty$. As mentioned in Hartman:2018tkw , Dirichlet boundary conditions might be problematic in higher dimensions in the sense of a well-posed initial valued problem Anderson_2008 ; An:2021fcq ; while the QFT is not guaranteed to exist; however, Hartman:2018tkw provided evidence in favor of (8), which we treat as our working assumption.
Following this line of work, we consider a sharp radial cutoff in global AdSd+1 space, with a geometry (9) for $N(r)=r^{2}+1$, where we can identify the coordinate change with respect to (7) for our choice $\ell_{d+1}=1$, $H=1$,
$r=\sinh\sigma\leavevmode\nobreak\ .$ (10)
Meanwhile, the $T^{2}$ deformed CFTd theory corresponding to an AdSd+1 bulk dual in Einstein gravity, is described by the action
$\partialderivative{I}{\lambda}=\int\mathrm{d}^{d}x\sqrt{h}\quantity[\quantity(T_{ij}+\frac{\alpha_{d}}{\lambda^{\frac{d-2}{d}}}G_{ij})^{2}-\frac{1}{d-1}\quantity(T_{i}^{i}+\frac{\alpha_{d}}{\lambda^{\frac{d-2}{d}}}G_{i}^{i})^{2}]\leavevmode\nobreak\ ,$ (11)
[…]
Using the coordinate change (10), we can identify the holographic dictionary relating the finite radial cutoff of AdSd+1 at $r=\sinh\sigma_{c}$, and the $T^{2}$ deformation parameter $\lambda$ as
$\lambda=\frac{4\pi G}{d\quantity(\sinh\sigma_{c})^{d}}\leavevmode\nobreak\ .$ (13)